# Ultra-Fast Language Generation via Discrete Diffusion Divergence Instruct

### v2

@@ Appendix C Student Objective
We begin by formally defining the objective $\mathcal{L}(\nu)$ as the Integral KL divergence. We assume the mask prior $\mathbf{m}$ has full support on the token simplex. Since the teacher uses the same absorbing kernel, $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$ share support (in particular, $\mathrm{supp}(\mathbf{q}_{\nu}(\mathbf{z}_{t},t))\subseteq\mathrm{supp}(\mathbf{q}_{\theta}(\mathbf{z}_{t},t))$), so $KL\left(\mathbf{q}_{\nu}(\mathbf{z}_{t},t)\,\big\|\,\mathbf{q}_{\theta}(\mathbf{z}_{t},t)\right)$ is well-defined and finite for all $t\in[0,1]$. Subsequently, we define the Objective $\mathcal{L}(\nu)$ as the Integral KL Divergence $D_{KL}(q,p)$ between the teacher and the student:
@@
Under mild regularity (bounded $\omega(t)$, dominated convergence, and differentiability of $\mathbf{p}_{\nu}(\mathbf{z}_{t},t)$ w.r.t. $\nu$), we can interchange $\nabla_{\theta}$ with the expectation and the time integral (and apply Fubini/Tonelli as needed). To evaluate the gradient of $\mathbb{E}_{\mathbf{z}_{t}\sim\mathbf{q}_{\nu}}\left[\log\mathbf{q}_{\nu}(\mathbf{z}_{t},t)-\log\mathbf{q}_{\theta}(\mathbf{z}_{t},t)\right]$ w.r.t. $\nu$, we use the identity $\nabla_{\theta}\mathbb{E}_{y\sim p_{\theta}}[f(y)]=\mathbb{E}_{y\sim p_{\theta}}[f(y)\nabla_{\theta}\log p_{\theta}(y)]+\mathbb{E}_{y\sim p_{\theta}}[\nabla_{\theta}f(y)]$. Applying this to our term gives:
@@
where $(i)$ applies the score-function (log-derivative) identity and is justified by moving $\nabla_{\theta}$ under the expectation. To derive $(ii)$, we know that $\nabla_{\nu}\log\mathbf{q}_{\theta}(\mathbf{z}_{t},t)=0$. Also, the following also holds true:
$\begin{split}\mathbb{E}_{\mathbf{z}_{t}\sim\mathbf{q}_{\nu}}\left[\nabla_{\nu}\log\mathbf{q}_{\nu}(\mathbf{z}_{t},t)\right]&=\sum_{\mathbf{z}_{t}}\mathbf{q}_{\nu}(\mathbf{z}_{t},t)\frac{\nabla_{\nu}\mathbf{q}_{\nu}(\mathbf{z}_{t},t)}{\mathbf{q}_{\nu}(\mathbf{z}_{t})}\\ &=\sum_{\mathbf{z}_{t}}\nabla_{\nu}\mathbf{q}_{\nu}(\mathbf{z}_{t},t)\\ &=\nabla_{\nu}\sum_{\mathbf{z}_{t}}\mathbf{q}_{\nu}(\mathbf{z}_{t},t)\\ &=\nabla_{\nu}(1)=0.\end{split}$
Subsequently, $(iii)$ rewrites $\nabla_{\nu}\log\mathbf{q}_{\nu}(\mathbf{z}_{t},t)$ as $\mathbb{E}_{\mathbf{x}\sim\mathbf{p}_{\nu}}[\nabla_{\nu}\log\mathbf{p}_{\nu}(\mathbf{z}_{t}=\mathbf{m},t=1)]$. We provide a detailed derivation here:
@@
where step $(i)$ uses $\mathbf{q}_{\nu}(\mathbf{z}_{t},t)=\mathbb{E}_{\mathbf{x}\sim\mathbf{p}_{\nu}}[\mathcal{Q}(\mathbf{z}_{t}\mid\mathbf{x})]$ and pass $\nabla_{\nu}$ through the expectation by linearity, step $(ii)$ applies $\nabla_{\nu}\mathbf{p}_{\nu}(\mathbf{z}_{t}=\mathbf{m},t=1)=\mathbf{p}_{\nu}(\mathbf{z}_{t}=\mathbf{m},t=1)\nabla_{\nu}\log\mathbf{p}_{\nu}(\mathbf{z}_{t}=\mathbf{m},t=1)$ to factor out a log-gradient, and step $(iii)$ consider Bayes rule $\mathbf{p}_{\nu}(\mathbf{z}_{t},t)=\mathbf{p}_{\nu}(\mathbf{x}\mid\mathbf{z}_{t})=\frac{\mathcal{Q}(\mathbf{z}_{t}\mid\mathbf{x})\mathbf{p}_{\nu}(\mathbf{z}_{t}=\mathbf{m},t=1)}{\mathbf{q}_{\nu}(\mathbf{z}_{t},t)}$.
@@
$\stackrel{{\scriptstyle}}{{=}}\int_{0}^{1}\omega(t)\mathbb{E}_{\mathbf{x}\sim\mathbf{p}_{\nu},\ \mathbf{z}_{t}\sim\mathcal{Q}}\left[R\left(\mathbf{z}_{t},t\right)\cdot\nabla_{\nu}\log\mathbf{p}_{\nu}\left(\mathbf{z}_{t}=\mathbf{m},t=1\right)\right]\mathrm{d}t,$
@@
Remark C.1.
In practice, this is approximated using a Monte Carlo estimator. The integral over time $t$ is replaced with an expectation over a sampling distribution $\pi(t)$, resulting in a single expectation over all random variables $(t,\mathbf{x},\mathbf{z}_{t})$. This unified expectation is then estimated by taking the sample mean over a mini-batch of size $N$:
$\nabla_{\nu}\mathcal{L}(\nu)\approx\frac{1}{N}\sum_{i=1}^{N}\left[\frac{\omega\left(t_{i}\right)}{\pi\left(t_{i}\right)}\cdot R\left(\mathbf{z}_{i},t_{i}\right)\cdot\nabla_{\nu}\log\mathbf{p}_{\nu}\left(\mathbf{z}_{t}=\mathbf{m},t=1\right)\right]$ (15)
@@
Appendix D Auxiliary Discriminator for Density Ratio Estimation
Our student objective, as derived in Appendix C, relies on the reward term $R(\mathbf{z}_{t},t):=\log\mathbf{q}_{\nu}(\mathbf{z}_{t},t)-\log\mathbf{q}_{\theta}(\mathbf{z}_{t},t)$, which is intractable due to the unknown marginal distributions $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$. To overcome this, we introduce a tractable estimator for the density ratio $\frac{\mathbf{q}_{\nu}(\mathbf{z}_{t})}{\mathbf{q}_{\theta}(\mathbf{z}_{t})}$ by training an auxiliary discriminator network. This approach is inspired by the principles of Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Wang et al., 2023) and has been successfully applied in Wang et al. (2025). We hereby adapt the following lemma to MDM setting.
For clarity, we adopt a time-agnostic notation where the corruption level is parameterized by the masking ratio $t\in[0,1]$, which corresponds to the schedule $\alpha_{t}$ via $t=1-\alpha_{t}$. We also denote $\mathbf{z}_{t}\in\mathcal{V}^{L}$ as a sequence with a mask ratio at time $t$.
Lemma D.1 (Density Ratio Representation, Restatement of Lemma 4.2).
Let $\mathbf{q}_{\theta}(\mathbf{z}_{t},t)$ and $\mathbf{q}_{\nu}(\mathbf{z}_{t},t)$ be the teacher and student model, respectively, over the discrete state space $\mathcal{V}^{L}$ at a mask ratio $\alpha_{t}$. Consider a discriminator $D_{\lambda}:\mathcal{V}^{L}\times[0,1]\to(0,1)^{L}$, which outputs a probability for each position $\ell\in\{1,\dots,L\}$. The discriminator $D_{\lambda}$ is trained to minimize the objective
@@ Appendix D Auxiliary Discriminator for Density Ratio Estimation
We here provide a detailed derivation for the optimal discriminator in the context of our discrete state space. The objective $\mathcal{L}_{D}(\lambda)$ is an expectation over the discrete random variable $\mathbf{z}_{t}$. We can express this expectation as a summation over all possible sequences $\mathbf{z}_{t}\in\mathcal{V}^{L}$:
$\mathcal{L}_{D}(\lambda)=\sum_{\mathbf{z}_{t}\in\mathcal{V}^{L}}\left[-\mathbf{q}_{\nu}(\mathbf{z}_{t},t)\log D_{\lambda}(\mathbf{z}_{t},t)-\mathbf{q}_{\theta}(\mathbf{z}_{t},t)\log(1-D_{\lambda}(\mathbf{z}_{t},t))\right]$
@@
$\frac{\partial}{\partial D_{\lambda}(\mathbf{z}_{t}^{\prime},t)}\mathcal{L}_{D}(\lambda)$ $=\frac{\partial}{\partial D_{\lambda}(\mathbf{z}_{t}^{\prime},t)}\left[-\mathbf{q}_{\nu}(\mathbf{z}_{t}^{\prime},t^{\prime})\log D_{\lambda}(\mathbf{z}_{t}^{\prime},t)-\mathbf{q}_{\theta}(\mathbf{z}_{t}^{\prime},t^{\prime})\log(1-D_{\lambda}(\mathbf{z}_{t}^{\prime},t))\right]$ (17)
$\stackrel{{\scriptstyle}}{{=}}-\frac{\mathbf{q}_{\nu}(\mathbf{z}_{t}^{\prime},t^{\prime})}{D_{\lambda}(\mathbf{z}_{t}^{\prime},t)}+\frac{\mathbf{q}_{\theta}(\mathbf{z}_{t}^{\prime},t^{\prime})}{1-D_{\lambda}(\mathbf{z}_{t}^{\prime},t)},$
which follows from standard differentiation of the logarithm function. We set (17) to be zero and get:
@@
Remark D.2.
The proof confirms that Lemma D.1 is robust to the specific properties of our MDM framework, including the discrete state space and the use of a mask ratio $t$ as the conditioning variable instead of a continuous time $t$. By training an auxiliary discriminator network $D_{\lambda}$ to approximate $D^{*}$, we obtain a tractable, principled estimator for the density ratio, which in turn provides a computable reward signal $\hat{R}(\mathbf{z}_{t},t)=\frac{1}{M}\sum_{\ell,\mathbf{z}_{t}^{\ell}=\mathbf{m}}\log\left(\frac{D_{\lambda}^{\ell}(\mathbf{z}_{t},t)}{1-D_{\lambda}^{\ell}(\mathbf{z}_{t},t)}\right)$ for our student objective, where $M$ denotes the number of masked tokens in the sequence.
Appendix E Additional Experimental Setup
This section provides a comprehensive overview of the configurations used in our experiments to ensure full reproducibility.
E.1 Model Architecture
@@ E.2 Dataset Details
E.3 Training and Distillation Hyperparameters
The detailed hyperparameters for teacher model pre-training and student model distillation are provided in Table 5. For our 169M models, we pre-trained the teacher model for 5 epochs. The subsequent distillation process was completed in about one hour on a single H100 GPU. For the 424M models, the teacher was pre-trained for 7 epochs, while the corresponding distillation took approximately 2 hours on a single H100 GPU.
E.4 Implementation Details
Our experiments were implemented using PyTorch. Key libraries and their versions are listed in Table 6. We plan to release our code and model checkpoints upon acceptance of the paper.
Table 6: Software and Library Versions.
Software Version
Python 3.12.11
PyTorch 2.5.1
Transformers 4.55.4
Datasets 4.0.0
Tokenizers 0.21.4
NumPy 2.3.2
SciPy 1.16.1
scikit-learn 1.7.1
Pandas 2.3.2
CUDA 12.1.0
cuDNN 9.1.0
Triton 3.1.0
E.5 Evaluation Details
Generative Perplexity: We use the implementation from the transformers library to compute the perplexity of generated samples under a frozen GPT-2 Large model. Text is processed identically to the training data.
@@ E.5 Evaluation Details
Appendix F Additional Experimental Results
F.1 Trade-Off Between Quality and Inference Speed
Figure 3: Perplexity versus latency trade-off, comparing RGAS against AS, the teacher model, and a GPT-2 baseline. The x-axis represents wall-clock latency in seconds per sequence (log scale), and the y-axis represents perplexity. Our method consistently achieves a superior efficiency frontier, reaching lower perplexity than AS at all latency points while approaching the quality of the teacher model with significantly less computational cost.
F.2 Results on Zero-Shot Perplexities

### v3

@@ Appendix B Student Objective Derivation
We begin by formally defining the objective $\mathcal{L}(\nu)$ as the Integral KL divergence. We assume the mask prior $\mathbf{m}$ has full support on the token simplex. Since the teacher uses the same absorbing kernel, $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$ share support (in particular, $\mathrm{supp}(\mathbf{q}_{\nu}(\mathbf{z}_{t},t))\subseteq\mathrm{supp}(\mathbf{q}_{\theta}(\mathbf{z}_{t},t))$), so $KL\left(\mathbf{q}_{\nu}(\mathbf{z}_{t},t)\,\big\|\,\mathbf{q}_{\theta}(\mathbf{z}_{t},t)\right)$ is well-defined and finite for all $t\in[0,1]$. We define the objective $\mathcal{L}(\nu)$ as the Integral KL Divergence between the teacher and the student:
@@
Under mild regularity conditions (bounded $\omega(t)$, dominated convergence, and differentiability of $\mathbf{p}_{\nu}(\mathbf{z}_{t},t)$ w.r.t. $\nu$), we can interchange $\nabla_{\theta}$ with the expectation and the time integral (and apply Fubini/Tonelli theorem as needed). To evaluate the gradient of $\mathbb{E}_{\mathbf{z}_{t}\sim\mathbf{q}_{\nu}}\left[\log\mathbf{q}_{\nu}(\mathbf{z}_{t},t)-\log\mathbf{q}_{\theta}(\mathbf{z}_{t},t)\right]$ w.r.t. $\nu$, we use the identity $\nabla_{\theta}\mathbb{E}_{y\sim p_{\theta}}[f(y)]=\mathbb{E}_{y\sim p_{\theta}}[f(y)\nabla_{\theta}\log p_{\theta}(y)]+\mathbb{E}_{y\sim p_{\theta}}[\nabla_{\theta}f(y)]$. Applying it to $\mathrm{T}_{1}$ yields
@@
where step $(i)$ applies the score-function (log-derivative) identity and is justified by moving $\nabla_{\theta}$ under the expectation. Step (ii) uses $\nabla_{\nu}\log\mathbf{q}_{\theta}(\mathbf{z}_{t},t)=0$ and the fact that
$\begin{split}\mathbb{E}_{\mathbf{z}_{t}\sim\mathbf{q}_{\nu}}\left[\nabla_{\nu}\log\mathbf{q}_{\nu}(\mathbf{z}_{t},t)\right]=\sum_{\mathbf{z}_{t}}\mathbf{q}_{\nu}(\mathbf{z}_{t},t)\frac{\nabla_{\nu}\mathbf{q}_{\nu}(\mathbf{z}_{t},t)}{\mathbf{q}_{\nu}(\mathbf{z}_{t})}=\sum_{\mathbf{z}_{t}}\nabla_{\nu}\mathbf{q}_{\nu}(\mathbf{z}_{t},t)&=\nabla_{\nu}\sum_{\mathbf{z}_{t}}\mathbf{q}_{\nu}(\mathbf{z}_{t},t)\\ &=\nabla_{\nu}(1)=0.\end{split}$
Step $(iii)$ rewrites $\nabla_{\nu}\log\mathbf{q}_{\nu}(\mathbf{z}_{t},t)$ as $\mathbb{E}_{\mathbf{x}\sim\mathbf{p}_{\nu}}[\nabla_{\nu}\log\mathbf{p}_{\nu}(\mathbf{z}_{t}=\mathbf{m},t=1)]$. A detailed derivation is as follows:
@@
where step $(i)$ uses $\mathbf{q}_{\nu}(\mathbf{z}_{t},t)=\mathbb{E}_{\mathbf{x}\sim\mathbf{p}_{\nu}}[\mathcal{Q}(\mathbf{z}_{t}\mid\mathbf{x})]$ and pass $\nabla_{\nu}$ through the expectation by linearity, step $(ii)$ applies $\nabla_{\nu}\mathbf{p}_{\nu}(\mathbf{z}_{t}=\mathbf{m},t=1)=\mathbf{p}_{\nu}(\mathbf{z}_{t}=\mathbf{m},t=1)\nabla_{\nu}\log\mathbf{p}_{\nu}(\mathbf{z}_{t}=\mathbf{m},t=1)$ to factor out a log-gradient, and step $(iii)$ follows from Bayes’ rule $\mathbf{p}_{\nu}(\mathbf{z}_{t},t)=\mathbf{p}_{\nu}(\mathbf{x}\mid\mathbf{z}_{t})=\frac{\mathcal{Q}(\mathbf{z}_{t}\mid\mathbf{x})\mathbf{p}_{\nu}(\mathbf{z}_{t}=\mathbf{m},t=1)}{\mathbf{q}_{\nu}(\mathbf{z}_{t},t)}$.
@@
$\stackrel{{\scriptstyle}}{{=}}\int_{0}^{1}\omega(t)\mathbb{E}_{\mathbf{x}\sim\mathbf{p}_{\nu},\ \mathbf{z}_{t}\sim\mathcal{Q}}\left[R\left(\mathbf{z}_{t},t\right)\cdot\nabla_{\nu}\log\mathbf{p}_{\nu}\left(\mathbf{z}_{t}=\mathbf{m},t=1\right)\right]\mathrm{d}t,$ (15)
@@
Remark B.1.
In practice, the objective in (15) is approximated using a Monte Carlo estimator. The integral over time $t$ is replaced with an expectation over a sampling distribution $\pi(t)$, resulting in a single expectation over all random variables $(t,\mathbf{x},\mathbf{z}_{t})$. This unified expectation is then estimated by taking the sample mean over a mini-batch of size $N$:
$\nabla_{\nu}\mathcal{L}(\nu)\approx\frac{1}{N}\sum_{i=1}^{N}\left[\frac{\omega\left(t_{i}\right)}{\pi\left(t_{i}\right)}\cdot R\left(\mathbf{z}_{t_{i}},t_{i}\right)\cdot\nabla_{\nu}\mathbf{p}_{\nu}\left(\mathbf{z}_{t}=\mathbf{m},t=1\right)\right]$ (16)
@@
Appendix C Auxiliary Discriminator for Density Ratio Estimation
Our student objective, as derived in Appendix B, relies on the reward term $R(\mathbf{z}_{t},t):=\log\mathbf{q}_{\nu}(\mathbf{z}_{t},t)-\log\mathbf{q}_{\theta}(\mathbf{z}_{t},t)$, which is intractable due to the unknown marginal distributions $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$. To overcome this, we introduce a tractable estimator for the density ratio $\frac{\mathbf{q}_{\nu}(\mathbf{z}_{t})}{\mathbf{q}_{\theta}(\mathbf{z}_{t})}$ by training an auxiliary discriminator network. This approach is inspired by the principles of Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Wang et al., 2023) and has been successfully applied in Wang et al. (2025). We hereby adapt the following lemma to MDM setting.
For clarity, we adopt a time-agnostic notation where the corruption level is parameterized by the masking ratio $t\in[0,1]$, which corresponds to the schedule $\alpha_{t}$ via $t=1-\alpha_{t}$. Let $\mathbf{z}_{t}\in\mathcal{V}^{L}$ denote a sequence with a mask ratio at time $t$.
Lemma C.1 (Density Ratio Representation).
Let $\mathbf{q}_{\theta}(\mathbf{z}_{t},t)$ and $\mathbf{q}_{\nu}(\mathbf{z}_{t},t)$ be the teacher and student models respectively, over the discrete state space $\mathcal{V}^{L}$ at a mask ratio $t$. Consider a discriminator $D_{\lambda}:\mathcal{V}^{L}\times[0,1]\to(0,1)^{L}$, which outputs a probability for each position $\ell\in\{1,\dots,L\}$. The discriminator $D_{\lambda}$ is trained to minimize the objective
@@ Appendix C Auxiliary Discriminator for Density Ratio Estimation
We here provide a detailed derivation for the optimal discriminator in the context of our discrete state space. The objective $\mathcal{J}(D)$ is an expectation over the discrete random variable $\mathbf{z}_{t}$. We can express this expectation as a summation over all possible sequences $\mathbf{z}_{t}\in\mathcal{V}^{L}$:
$\mathcal{J}(D)=\sum_{\mathbf{z}_{t}\in\mathcal{V}^{L}}\left[-\mathbf{q}_{\nu}(\mathbf{z}_{t},t)\log D_{\lambda}(\mathbf{z}_{t},t)-\mathbf{q}_{\theta}(\mathbf{z}_{t},t)\log(1-D_{\lambda}(\mathbf{z}_{t},t))\right]$
@@
$\frac{\partial}{\partial D_{\lambda}(\mathbf{z}_{t}^{\prime},t)}$ $\left[-\mathbf{q}_{\nu}(\mathbf{z}_{t}^{\prime},t^{\prime})\log D_{\lambda}(\mathbf{z}_{t}^{\prime},t)-\mathbf{q}_{\theta}(\mathbf{z}_{t}^{\prime},t^{\prime})\log(1-D_{\lambda}(\mathbf{z}_{t}^{\prime},t))\right]$ (18)
$\stackrel{{\scriptstyle(i)}}{{=}}-\frac{\mathbf{q}_{\nu}(\mathbf{z}_{t}^{\prime},t^{\prime})}{D_{\lambda}(\mathbf{z}_{t}^{\prime},t)}+\frac{\mathbf{q}_{\theta}(\mathbf{z}_{t}^{\prime},t^{\prime})}{1-D_{\lambda}(\mathbf{z}_{t}^{\prime},t)}$
where step $(i)$ follows from standard differentiation of the logarithm function. We set (18) to be zero and get:
@@
Remark C.2.
The proof confirms that Lemma C.1 is robust to the specific properties of our MDM framework, including the discrete state space and the use of a mask ratio $t$ as the conditioning variable instead of a continuous time $t$. By training an auxiliary discriminator network $D_{\lambda}$ to approximate $D^{*}$, we obtain a tractable, principled estimator for the density ratio, which in turn provides a computable reward signal
$\hat{R}(\mathbf{z}_{t},t)=\frac{1}{M}\sum_{\ell,\mathbf{z}_{t}^{\ell}=\mathbf{m}}\log\left(\frac{D_{\lambda}^{\ell}(\mathbf{z}_{t},t)}{1-D_{\lambda}^{\ell}(\mathbf{z}_{t},t)}\right)$
for our student objective, where $M$ denotes the number of masked tokens in the sequence.
Appendix D Additional Experimental Setup
D.1 Baseline Setup
To ensure a fair and reproducible comparison, we strictly follow the official implementations and experimental configurations of both SDTT and DUO. For SDTT (Deschenaux and Gulcehre, 2025), we adopt the same 169M MDM teacher used in our method, pretrained for 1024 NFEs, and perform seven rounds of distillation with 10K steps each (70K steps in total) to obtain an 8-step student, exactly matching the protocol in the official repository. For DUO (Sahoo et al., 2025), we use the 169M USDM teacher (1024 NFEs) released in their codebase and conduct five rounds of distillation (50K total steps). We additionally evaluate their pretrained few-step student model included in the repository.
In practice, we observe non-negligible variation in the performance of these baselines across different training runs. To provide the most stable and equitable comparison, we report the metrics published in their original papers, which also match the strongest results we reproduced internally.
D.2 Model Architecture
@@ D.3 Dataset Details
D.4 Training and Distillation Hyperparameters
The detailed hyperparameters for teacher model pre-training and student model distillation are provided in Table 5. For our 169M models, we pre-trained the teacher model for around 200,000 steps. The subsequent distillation process was completed in about one hour on a single H100 GPU. For the 424M models, the teacher was pre-trained for around 350,000 steps, while the corresponding distillation took approximately two hours on a single H100 GPU.
D.5 Evaluation Details
Generative perplexity: We use the implementation from Hugging Face’s transformers library to compute the perplexity of generated samples under a frozen GPT-2 Large model. Text is processed identically to the training data.
@@ D.5 Evaluation Details
Appendix E Additional Experimental Results
E.1 Diversity and Fidelity Analysis
To provide a comprehensive comparison beyond perplexity, we evaluate sample diversity using MAUVE (for distribution similarity) and Self-BLEU (for diversity). All models generate 1,024 unconditional samples with identical decoding configurations (no prompt, max length 1024). Baselines use the same sampling budgets (8, 16, 32, 64, and 128 NFEs), and all samples are processed using the same GPT-2 tokenizer.
MAUVE measures the similarity between the generated distribution $Q$ and the reference distribution $P$ by analyzing the trade-off between Type I and Type II errors. Let $R_{\lambda}=\lambda P+(1-\lambda)Q$ be a mixture distribution for $\lambda\in(0,1)$. We compute the divergence frontier as a curve $(x(\lambda),y(\lambda))$ in the 2D plane, where the coordinates are exponentially scaled Kullback-Leibler (KL) divergences:
$x(\lambda)=\exp\left(-c\cdot D_{\mathrm{KL}}(Q\|R_{\lambda})\right)\quad\text{and}\quad y(\lambda)=\exp\left(-c\cdot D_{\mathrm{KL}}(P\|R_{\lambda})\right),$ (20)
where $c>0$ is a scaling factor (we set $c=5$). As $\lambda$ varies, these coordinates trace a curve within the unit square $0,1]^{2}$. The MAUVE score is calculated as the area under this curve using numerical integration (trapezoidal rule) over a discretized set of $\lambda$ values. Normally, texts are truncated to 256 tokens before embedding.
Self-BLEU evaluates diversity through n-gram overlap. For $N$ samples $\{\mathbf{x}_{i}\}_{i=1}^{N}$:
$\text{Self-BLEU}=\frac{1}{N}\sum_{i=1}^{N}\mathrm{BLEU}(\mathbf{x}_{i},\{\mathbf{x}_{j}\}_{j\neq i}),\qquad\text{where }\mathrm{BLEU}=BP\cdot\exp\left(\sum_{n=1}^{5}w_{n}\log p_{n}\right),$
where $p_{n}$ is the n-gram precision, $w_{n}=0.2$, and $BP$ is the brevity penalty. We set $n=5$, use the smoothing method 1, and the tokenization of NLTK.
Table 6: Additional generation metrics: MAUVE ($\uparrow$, higher is better) and Self-BLEU ($\downarrow$, lower indicates more diversity). Comparison across varying NFEs.
Metrics Methods 8 NFEs 16 NFEs 32 NFEs 64 NFEs 128 NFEs
MAUVE ($\uparrow$) MDLM 4.77 6.76 6.46 5.50 5.97
($\times 10^{-3}$) SDTT 5.60 5.38 5.70 5.04 5.96
DUO 5.75 5.15 5.99 5.85 6.05
DiDi-Instruct (Ours) 5.84 6.55 6.54 6.14 6.31
Self-BLEU ($\downarrow$) MDLM 2.91 4.58 6.36 7.97 9.33
($\times 10^{-2}$) SDTT 5.02 7.98 10.09 12.49 12.92
DUO 8.26 9.28 9.62 8.84 8.94
DiDi-Instruct (Ours) 4.82 4.64 5.24 6.25 8.24
Table 6 presents the full results. DiDi-Instruct consistently achieves the highest MAUVE scores across all NFEs, which validates that our adversarial distribution-matching objective successfully aligns the student’s marginals with the teacher’s. Self-BLEU values remain competitive, demonstrating that diversity is not sacrificed for fidelity. Intuitively, baselines such as SDTT and DUO tend to prioritize pointwise accuracy, leading to mode-seeking behavior and worse Self-BLEU. In contrast, score decomposition forces the student to model noisy state evolution, preserving diverse modes, while adversarial training ensures marginal distribution matching for higher fidelity.
E.2 Practical Latency and Throughput Analysis
To complement the NFE-based analysis in the main text, we conduct a practical latency evaluation. All experiments use a single H100 GPU with bfloat16 precision, batch size $16$, and sequence length $1024$. All models share the same architecture as demonstrated in Table 3. The AR baseline uses greedy decoding with standard optimizations such as KV-caching and FlashAttention.
Table 7: Practical latency and throughput comparison on one H100 GPU.
Models NFEs Tokens/sec Latency per 1K tokens (sec)
AR 1024 179.042 5.719
MDLM 512 111.146 9.213
SDTT 64 806.498 1.269
DiDi-Instruct (Ours) 16 2366.604 0.432
The latency across different methods is reported in Table 7. At matched perplexity, DiDi-Instruct achieves a $13.2\times$ reduction in latency per 1k tokens relative to AR, while maintaining high sample quality. This confirms that the advantages of distilled diffusion models persist under realistic decoding settings and not merely under NFE-based comparisons. Note that the measured latency improvement is a conservative estimate: DiDi-Instruct is fully compatible with hybrid AR–diffusion architectures (Arriola et al., 2025), and benefits from recent advances enabling KV-cache reuse and parallel diffusion decoding (Wu et al., 2026), which offer additional opportunities for further acceleration.
E.3 Results on zero-shot perplexities
As detailed in Table 8, DiDi-Instruct demonstrates superior generalization capabilities compared to the distilled baseline. It achieves lower perplexity across different datasets relative to DUO distilled. On benchmarks like Wikitext, it notably surpasses the autoregressive GPT-2 baseline. These results confirm that DiDi-Instruct effectively preserves the teacher’s semantic coverage, mitigating the mode collapse typically associated with few-step diffusion generation.